OpenStudy (anonymous):

(x-3)^2/(9)-(y-2)^2/(16)=1 find the center of the hyperbola

OpenStudy (anonymous):

\[(x-3)^2/(9)-(y-2)^2/(16)=1\]

OpenStudy (anonymous):

Which is ( h, k)!

sam (.sam.):

Set each brackets =0 (x-3)=0 x=3 (y-2)=0 y=2